OpenStudy (anonymous):

Find the rate of change y=-3x+1

OpenStudy (pradipgr817):

take dy/dx, it is -3

Parth (parthkohli):

Our rate of change is simply the \(\text{slope}\) in this case.

Parth (parthkohli):

@jmallis Do you not agree with this? We just have a line!

Parth (parthkohli):

Why does one need to take stuff to calculus?

OpenStudy (pradipgr817):

yes slope is constant

OpenStudy (across):

@ParthKohli, that's because they don't really understand calculus. ;P

Parth (parthkohli):

Yes. In a linear equation, rate of change is simply the slope... right?

OpenStudy (pradipgr817):

when you can use it. you use it

OpenStudy (across):

@pradipgr817, I usually slice my chicken with a chainsaw.

Parth (parthkohli):

What if a kid is in the sixth grade? Will you really care to explain what \(d \over dx\) is? Just ask him what number he sees along with \(x\)!

OpenStudy (pradipgr817):

thats your choice

Parth (parthkohli):

So, very, ridiculous.

OpenStudy (pradipgr817):

@ParthKohli thats you know he's in sixth. you went you sent simple.

OpenStudy (across):

Let's all follow @@pradipgr817's school of thought! @jmallis, to find the slope of the super complicated function\[y=-3x+1,\]you have to take its derivative, as follows:\[\frac{dy}{dx}=\lim_{\Delta x\to0}\frac{-3(x+\Delta x)+1+3x-1}{\Delta x}=\lim_{\Delta x\to0}\frac{-3\Delta x}{\Delta x}=-3.\]Did you get that? :P

OpenStudy (anonymous):

Ok

OpenStudy (pradipgr817):

@ across, you got it right

Parth (parthkohli):

She knows much more Mathematics than you, @pradipgr817

OpenStudy (pradipgr817):

@ I told what came instantly after looking the problem.